Suppose you deposit $6000 at 7% interest compounded continuously. Find the average value of your account during the first 3 years.
Answer #1

A = pert 0.07t A=6000 3 Average valuez 6000 60077 dt 3-0 E = feroj 004 C e907t] [eootuse Σ 2000 0.07 2000 0.07 = 6676-5159987
Answer #2

SOLUTION :


FOR CONTINUOUS COMPOUNDING :


FV = A = A0 e^(r t) = 6000 * e^(0.07 * 3) = 7402.07 ($) 


Average value of the account during first 2 years 

= (A0 + A) / 2 

= (6000 + 7402.07)/2 

= 6701.04 ($) (ANSWER).

answered by: Tulsiram Garg

> 3rd line : ......." first 2 year"s may please be corrected as "first 3 years".

Tulsiram Garg Wed, Jan 19, 2022 7:37 AM